Operations Report – May 04th

SOL: 12

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10.

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 8:45 pm to 7:25 am.

Solar— SOC 60 % (Before generator is run at night)

Diesel – 60 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 30 gallons.

Water (static) – 285 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134635.2

Toilet tank emptied: no
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 350.1, 350.3

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 2.5

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat. 300 would not start, there is a grinding or clicking sound, as if teeth are broken off of the starter or flywheel.

Deimos rover used: Yes
Hours: 125.6
Beginning charge: 100%
Ending charge: 73 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.8
Beginning charge: 100 %
Ending charge: 86 %
Currently charging: Yes

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: Still work to do on two piece suits (straps). Suit # 7 failed during the final EVA today; fans got stuck in shroud and released magic smoke.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment concluding.

Summary of RAM operations: Outreach video created in RAM today.

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Our garbage is ready to be picked up (both burnable and non burnable). We would also like more ATV fuel (ethanol free gasoline).

Copyright © The Mars Society. All rights reserved. | Main Site